# IberFire - a detailed creation of a spatio-temporal dataset for wildfire risk assessment in Spain

### v1

@@ 4 Technical Validation
Four distinct validation procedures were conducted to ensure the robustness and reliability of IberFire. First, the integrity of data formats and measurement units was verified to guarantee internal consistency across all features. Second, given that vegetation indices were derived from satellite-based remote sensing data, the presence of missing values was common. To address this, multiple imputation techniques were evaluated, and the method yielding the lowest reconstruction error was applied to fill missing data. Third, available AEMET historical measurement data were downloaded and compared with IberFire meteorological data. Finally, an XGBoost model was trained on data from 2008 to 2023 and fire risk maps were plotted in various 2024 days, which served as a practical validation step to assess the predictive utility of IberFire. This chapter provides a detailed account of each of these four validation strategies aimed at guaranteeing the quality and usability of the IberFire datacube.
@@ 4.1 Data correctness
A datacube is a multi-dimensional data structure that integrates spatial and temporal information in a unified framework. Within the IberFire datacube, this includes coordinate grids, binary and categorical attributes, proportion-based data, and continuous numerical features. Given the complexity and heterogeneity of these components, quality assurance procedures are needed to ensure the reliability and internal consistency of the datacube. This section outlines the validation processes carried out to assess the correctness of temporal coordinates, the integrity of categorical and binary variables, the validity of proportion-based attributes, and the credibility of numerical values. These checks were designed to detect anomalies and confirm that the data adheres to defined structural standard across the entire temporal range of the datacube.
First, a temporal consistency check was conducted to guarantee that there are no duplicated time coordinates in the datacube. The dataset was verified to contain valid accessible values for all spatio-temporal features across the entire temporal range from 01/12/2007 to 31/12/2024. For the spatial features, visual inspection was performed to ensure that spatial coordinates are homogeneous across all the region of interest.
For categorical and binary features, it was ensured that all entries conformed to the expected set of values. For the AutonomousCommunities feature, all labels were verified to fall within the valid range of 0 to 19, with no occurrences of invalid or unexpected values. Regarding binary features, is_spain, is_sea, is_waterbody, is_holiday, is_natura2000, is_fire, and is_near_fire, it was confirmed that all values were strictly limited to either 0 or 1.
@@
Additionally, for the proportion-based features derived from the CLC dataset, structured hierarchically into three levels as outlined in Section 2.5, consistency checks were performed to ensure that the proportions at each hierarchical level sum to one for every cell. Therefore, it was verified that the five aggregated features at Label 1, the fourteen intermediate-level features at Label 2, and the forty-four fine-grained features at Label 3 each sum to exactly one per cell.
Finally, for numerical features such as meteorological variables and vegetation indices, visual inspection was carried out through exploratory plotting to identify potential outliers and ensure that all values fell within logical and expected ranges. This process revealed a substantial number of missing values in certain data sources used to construct the vegetation indices. For instance, in the case of FAPAR, no missing values were observed until 30/04/2020, which corresponds to the ending of the first data source. After that date, the second source was used and exhibited a considerable number of gaps. The methodology adopted to address this issue is detailed in the following section.
@@ 4.2 Missing values validation
Once all features were validated to fall within their expected formats and value ranges, the missing values identified in the vegetation indices were subsequently addressed. To do so, different interpolation techniques were compared, and the one with the lowest reconstruction error was applied. This step was crucial, as not all machine learning algorithms can inherently manage missing data.
The imputation process was applied exclusively to the vegetation indices, as the remaining data sources were either complete or contained a negligible amount of missing values. In particular, no missing data were observed in any of the spatial-only features. Regarding spatio-temporal features, the layers is_fire, is_near_fire, and is_holiday were complete for the entire temporal range. Likewise, the meteorological variables derived from the ERA5-Land dataset did not exhibit missing values within the Spanish territory (is_spain = 1), due to the spatial continuity and post-processing of the source data.
To address the missing values in the vegetation indices, several imputation techniques were evaluated by introducing artificial gaps in dates known to be complete. The introduced gaps followed the real shapes of the missing values observed in satellite-derived vegetation indices. Subsequently, various imputation methods were applied to reconstruct the artificially masked values, and the reconstruction accuracy of each method was quantified using the Mean Absolute Error (MAE). This evaluation was performed independently for each vegetation index.
Four imputation techniques were selected for evaluation: nearest neighbour, linear, quadratic, and cubic interpolation. Each method was applied along the temporal axis, and linear interpolation yielded the lowest reconstruction errors across every tested feature. Consequently, it was adopted as the imputation method for filling missing values in the vegetation indices.
@@ 4.3 Data comparison with AEMET
This validation involved comparing the final, transformed ERA5-Land variables included in IberFire against historical records from AEMET meteorological stations. Particular attention was paid to the u-wind and v-wind components, which underwent the most complex transformations, as they required the reconstruction of wind speed and direction.
To perform the comparison, historical AEMET data were retrieved from
https://datosclima.es/Aemethistorico/Descargahistorico.html. This source provides daily records from meteorological stations, including maximum, mean, and minimum temperature, precipitation, mean wind speed, wind direction at maximum wind speed, maximum wind speed, and both maximum and minimum surface pressure. Unfortunately, this source does not include historical records for other meteorological variables featured in IberFire, such as relative humidity. No alternative source was found that offers historical records for these variables. Furthermore, the retrieved AEMET dataset contained missing values for the available variables, and many of those features were not provided in all stations.
For each meteorological station within the region of interest, the Mean Absolute Error (MAE) was calculated for all available dates between 01/01/2007 and 31/12/2024, across all available features. Selected examples of these results are shown in the top four images of Figure 6. As illustrated, precipitation records are available for a considerably larger number of stations as opposed to those of surface pressure. Not only that, some stations have higher MAE values than others on the same feature.
@@
[…]
@@
#NA_values 58 58 58 67 169 169 169 605 605
Table 9: Summary of MAE and NMAE metric values. Feature names from left to right: t2m_max, t2m_mean, t2m_min, total_precipitation_mean, wind_speed_mean, wind_direction_at_max_speed, wind_speed_max, surface_pressure_max, and surface_pressure_min.
The results indicate that the temperature features are highly reliable. Even in the worst-performing station, the mean Mean Absolute Error (MAE) remains as low as $2.173$, with a standard deviation of $1.215$. Assuming a normal distribution of errors, this implies that in 95% of the cases, the discrepancy between AEMET measurements and IberFire data does not exceed 4°. Given that temperature values at a single station can vary by up to 40°across different days, this error value is relatively small and supports the robustness of the temperature variables in the IberFire datacube. Moreover, as the temperature values are derived from ERA5-Land, a dataset rigorously validated by the Copernicus Climate Data Store, their reliability is further reinforced.
In comparison, the precipitation and surface pressure features exhibit higher NMAE values than temperature. Nevertheless, these variables are also sourced from Copernicus ERA5-Land, a dataset extensively validated. Consequently, despite their relatively higher error metrics, the integrity and reliability of these features remain supported by the quality and consistency of the underlying data source.
@@ 4.4 Fire risk mapping validation
To evaluate the practical applicability of the IberFire datacube in real-world scenarios, a fire risk mapping exercise was conducted. The objective was to assess whether the datacube features can support the generation of reliable fire risk maps when used in combination with standard machine learning techniques.
@@
As discussed in Chapter 5, forest fires are intrinsically rare events, resulting in a highly imbalanced distribution of the is_fire feature, with significantly fewer fire instances relative to non-fire occurrences. To address this imbalance during model training, a balanced dataset was retrieved from IberFire by including all fire instances recorded between 2008 and 2023, complemented by an approximately equal number of randomly selected non-fire instances from the same dates. Due to the stochastic nature of the non-fire instance sampling, the final class counts are not exactly equal. The resulting training dataset consisted of 140,399 instances, comprising 70,476 fire cases and 69,923 non-fire cases. For each selected instance, corresponding to a unique spatio-temporal cell, all associated features were extracted from the IberFire datacube and stored in CSV format, with each row representing a single cell and its corresponding attributes.
@@
Following the construction of the dataset, an XGBoost classifier was trained on the extracted instances. To reduce the risk of overfitting and ensure the generalisation capability of the model, a cross-validation strategy was implemented. After model development, an independent test set was constructed by retrieving a new balanced dataset from the IberFire datacube, consisting of all available fire instances from the year 2024 along with a comparable number of randomly sampled non-fire instances from the same dates. The trained model was then evaluated on this test set, yielding an Area Under the Curve (AUC) of 0.955, which indicates a high degree of predictive accuracy and robustness in real-world forecasting scenarios.
@@
In addition to this quantitative evaluation, the classifier was used to generate daily fire risk predictions for the entire year of 2024. For each day, all corresponding spatio-temporal instances were extracted from the IberFire datacube and converted into CSV format to serve as input for the model. Predictions were then computed for each instance. Subsequently, the auxiliary features x_index and y_index were used to construct the daily fire risk raster maps. A selection of these predicted maps is presented in Figure 7.
Figure 7: Forest fire risk maps for various dates of 2024.
@@
Visual inspection of the resulting fire risk maps revealed a high degree of spatial and temporal coherence and alignment with historical fire incidence patterns illustrated in Figure 1. Notably, areas with historically elevated fire activity, such as Galicia, Asturias, and regions along the Portuguese border, consistently exhibited higher predicted risk levels, reinforcing the model’s ability to capture meaningful spatial trends in fire susceptibility. Furthermore, the model demonstrated the capability to reflect temporal variability in fire risk, with elevated predictions on dates corresponding to confirmed fire events.
@@ 5 Usage notes
The IberFire datacube was developed to support the modelling of wildfire occurrence risk across the Spanish territory, excluding the Canary Islands. A spatio-temporal structure was implemented to fulfil this objective, providing daily data from 01/12/2007 to 31/12/2024 over a regular grid of 1km $\times$ 1km $\times$ spatial resolution. This design enables the training and validation of ML and DL models for wildfire risk prediction. The datacube is intended exclusively for mainland Spain and the Balearic Islands. To ensure that only relevant cells with complete and valid data are included in modelling workflows, users should filter the dataset accordingly using the variable is_spain (by selecting the instances where is_spain = 1).
@@
The datacube includes daily records beginning in December 2007; however, fire occurrence data were retrieved only from 01/01/2008 onward, and all cells for December 2007 were explicitly set as non-fire. The data for December 2007 is included to enable the modelling of DL models that need multiple days of past data, like long short-term memory (LSTM) networks. Thus, even though fire predictions focus on the period from 2008 to 2024, the extended time frame ensures that models requiring temporal dependencies can be effectively trained.
@@
Prediction targets can be defined using either the is_fire variable, which indicates whether a fire occurred in a given cell on a specific date, or the is_near_fire variable, which flags cells located in proximity to a fire event. The remaining features in the datacube are suitable for use as explanatory variables in predictive models. However, the use of the auxiliary variables x_index and y_index as model inputs is not recommended. These features were introduced to facilitate data filtering and descriptive analysis, but they lack geospatial meaning beyond the cell grid. Their values are specific to the internal indexing of the datacube and are not transferable to other geographic regions. For spatial modelling purposes, the variables x_coordinate and y_coordinate, which represent the actual projected coordinates of each grid cell on the EPSG:3035 CRS, can be used instead as they retain geographic meaning and are suitable for models requiring location-aware features.
@@
To incorporate information about the Spanish territorial division, it is recommended to apply one-hot encoding to the categorical AutonomousCommunities feature, resulting in 17 binary features corresponding to the autonomous communities listed in Table 1, excluding the Canary Islands, Ceuta, and Melilla, which are not part of the study area. Similarly, a month feature can be derived from the temporal coordinate of each spatio-temporal cell. As different months exhibit distinct seasonal fire risk patterns, it is also advisable to apply one-hot encoding to this variable.
@@
It is important to note that some features, such as those derived from CLC and population density (popdens), are not static but are also not temporally continuous. These variables are associated with a specific year of update and thus are included in the datacube with spatial dimensions only $(x,y)$, rather than full spatio-temporal indices $(x,y,t)$. When preprocessing the datacube for modelling applications or conversion to tabular formats such as CSV or DataFrame structures, it is crucial to ensure that for each spatio-temporal cell $(x,y,t)$, only information available at the time $t$ of the cell is used. For instance, when modelling a record for July 15, 2010, the appropriate population density value to use is popdens_2010, which should be assigned to a generic popdens column. Similarly, for a CLC-derived feature such as CLC_urban_fabric_proportion, the correct value to assign would be from CLC_2006_urban_fabric_proportion, since the next update in 2012 would not yet be available at that date. Careful consideration is required to prevent potential data leakage, particularly in the case of land cover features that might implicitly reflect fire occurrence. The CLC dataset contains a class representing area that was burned, which, if taken from a dataset version updated after the date being modelled, could inadvertently introduce information correlated with the target variable (is_fire). Such unintended inclusion of future knowledge may affect the integrity and validity of predictive modelling outcomes. When properly preprocessed, the resulting CSV dataset extracted from IberFire should contain a total of 145 columns, including one-hot encoded representations of month and autonomous communities.
@@
An essential final consideration when working with the IberFire datacube is the substantial computational resources required for data processing. Although the total disk size of the datacube is approximately 29GB, each spatio-temporal feature is stored in a highly compressed format. When decompressed, a single feature represented as a float32 array occupies around 25GB of memory, making it impractical to load multiple features into RAM simultaneously on standard computing hardware. The xarray package handles this problem by chunking the data, and for small computations like data visualisation and analysis, standard hardware suffices. Similarly, retrieving a balanced CSV training dataset, including all fire instances and an equal number of non-fire instances, is believed to be feasible under standard memory constraints, although this has not been empirically verified on standard hardware. However, more demanding operations, such as the generation of new spatio-temporal variables derived from combinations of existing features, require significantly more memory. Therefore, all the processing steps involved in the construction of IberFire, including preprocessing and training dataset extraction, were executed on a machine equipped with 128GB of RAM.

### v2

@@ 4 Technical Validation
Four distinct validation procedures were conducted to ensure the robustness and reliability of IberFire. First, the integrity of data formats and measurement units was verified to guarantee internal consistency across all features. Second, given that vegetation indices were derived from satellite-based remote sensing data, the presence of missing values was common. To address this, multiple imputation techniques were evaluated, and the method yielding the lowest reconstruction error was applied to fill missing data. Third, available AEMET historical measurement data were downloaded to compare with IberFire meteorological data. Finally, an XGBoost model was trained on data from 2008 to 2023 and using the predicted class probabilities, fire risk maps were plotted for various 2024 days, which served as a practical validation step. This section provides a detailed account of each of these four validation strategies aimed at guaranteeing the quality and usability of the IberFire datacube.
@@ 4.1 Data correctness
A datacube is a multi-dimensional data structure that integrates spatial and temporal information in a unified framework. Within the IberFire datacube, this includes grid coordinate values, binary and categorical attributes, proportion-based data, and continuous numerical features. Given the complexity and heterogeneity of these components, quality assurance procedures are needed to ensure the reliability and internal consistency of the datacube.
First, a temporal consistency check was conducted to guarantee that there are no duplicated time coordinates in the datacube. The dataset was verified to contain valid accessible values for all spatio-temporal features across the entire temporal range from 01/12/2007 to 31/12/2024. For the spatial features, visual inspection was performed to ensure that spatial coordinates are homogeneous across the entire region of interest.
For categorical and binary features, it was ensured that all entries conformed to the expected set of values. For the AutonomousCommunities feature, all labels were verified to fall within the valid integer range of 0 to 19, with no occurrences of invalid or unexpected values. Regarding binary features, is_spain, is_sea, is_waterbody, is_holiday, is_natura2000, is_fire, and is_near_fire, it was confirmed that all values were strictly limited to either 0 or 1.
@@
Additionally, for the proportion-based features derived from the CLC dataset, structured hierarchically into three levels described in Table 3, consistency checks were performed to ensure that the proportions at each hierarchical level sum to one for every cell. Therefore, it was verified that the five aggregated features at Label 1, the fourteen intermediate-level features at Label 2, and the forty-four fine-grained features at Label 3 each sum to exactly one per cell.
Finally, for numerical features such as meteorological variables and vegetation indices, visual inspection was carried out through exploratory plotting to identify potential outliers and ensure that all values fell within logical and expected ranges. This process revealed a substantial number of missing values in certain data sources used to construct the vegetation indices. For instance, in the case of FAPAR, no missing values were observed until 30/04/2020, which corresponds to the ending of the first data source [FAPAR_v1]. After that date, the second source was used [FAPAR_v2] and exhibited a considerable number of gaps. The methodology adopted to address this issue is detailed in the following subsection.
@@ 4.2 Missing values validation
Once all features were validated to fall within their expected formats and value ranges, the missing values identified in the vegetation indices were addressed. This step was crucial, as not all machine learning algorithms can inherently manage missing data.
The imputation process was applied exclusively to the vegetation indices, since all other data sources were either complete or contained a negligible amount of missing values. In particular, no missing data were observed in any of the spatial-only features, and regarding spatio-temporal features, the layers is_fire, is_near_fire, and is_holiday were complete for the entire temporal range. Likewise, the meteorological variables derived from the ERA5-Land dataset did not exhibit missing values within the Spanish territory (is_spain = 1), due to the spatial continuity and post-processing of the source data.
Four interpolation techniques were analysed to address the missing values in the vegetation indices: nearest neighbour, linear, quadratic, and cubic. Each method was applied along the temporal axis independently on each spatial cell.
To evaluate these methods, artificial gaps were first added in dates known to be complete. These gaps followed the real shapes of the missing values observed in satellite-derived vegetation indices. Specifically, 140 NAN masks were obtained from the FAPAR variable and inserted on all the indices. Three examples of these artificial masks are shown in the top image of Figure 6.
Figure 6: Top: Three examples of binary NAN masks that were introduced as artificial NAN values. Bottom: Comparison between original FAPAR values, masked FAPAR values and interpolated FAPAR values.
Subsequently, the four imputation methods were applied to reconstruct the artificially masked values. To evaluate the reconstruction accuracy, the absolute differences between the true and the imputed values were computed for each masked day. These differences were then summed across all data points for that day, and the resulting sums were averaged over all masked days (i.e., they were divided by 140). This yielded the Reconstruction Error (RE), which served as the primary metric for comparing the performance of the imputation methods.
The evaluation was performed independently for each vegetation index. To illustrate the effect of the interpolation, the bottom image in Figure 6 displays an example with the original data of FAPAR, the artificially masked version, and the reconstruction obtained using linear interpolation.
Table 10 presents the RE scores for each vegetation index across the four imputation methods. As shown in the table, linear interpolation yielded the lowest RE across every tested feature. Consequently, it was adopted as the imputation method for filling the real missing values in the vegetation indices. It should be noted that the RE varies a lot across different vegetation indices due to the difference in magnitudes, not because some indices have inherently more missing data or are harder to reconstruct. For example the NDVI ranges between -1 and 1 whereas the SWI ranges from 0 to 100. Furthermore, RE values are not expected to match the magnitude as the original feature values, since the RE represents the sum of absolute errors across all cells for a given day.
Tested variable Nearest neighbour Linear Quadratic Cubic
FAPAR 3412 2418 4089 4191
LAI 14705 9941 14716 15037
NDVI 4239 3588 * *
SWI_001 724876 649175 1507907 1601710
SWI_005 326488 273457 585050 608277
SWI_010 224292 177245 367986 381461
SWI_020 151797 112792 214688 222810
LST 325655 285057 * *
Table 10: Reconstruction Error (RE) of the interpolation techniques for all the tested variables
(* The interpolation technique was not assessed for that feature due to execution errors caused from characteristics of the data).
@@ 4.3 Data comparison with AEMET
This validation compared the transformed ERA5-Land variables included in IberFire against historical records from AEMET meteorological stations. Particular attention was paid to the u-wind and v-wind components, which underwent the most complex transformations, as they required the reconstruction of wind speed and direction.
To perform the comparison, historical AEMET data were retrieved from https://datosclima.es/Aemethistorico/Descargahistorico.html. This source provides daily records from meteorological stations, including maximum, mean, and minimum temperature; precipitation; mean wind speed; wind direction at maximum wind speed; maximum wind speed; and both maximum and minimum surface pressure. Unfortunately, this source does not include historical records for other meteorological variables featured in IberFire, such as relative humidity. No alternative source was found that offers historical records for these variables. Furthermore, the retrieved AEMET dataset contained missing values for the available variables, and many of those features were not provided in all stations.
For each meteorological station within the region of interest, the Mean Absolute Error (MAE) was calculated for all available dates between 01/01/2007 and 31/12/2024, for each available feature. To calculate it, the values of the nearest cell to each meteorological station were selected. Chosen examples of these results are shown in the top four images of Figure 7. As illustrated, precipitation records are available for a considerably larger number of stations compared to surface pressure. Additionally, some stations have higher MAE values than others for the same feature.
@@
$\text{NMAE(feature, station)}=\frac{\text{MAE(feature, station)}}{max(\text{feature, station})-min(\text{feature, station})}\,$, (5)
where MAE(feature, station) is the MAE value of a feature in a given station, max(feature, station)max($\text${feature, station}) is the maximum value of the feature measured in that station, and min(feature, station)min($\text${feature, station}) is the minimum. Table 11 provides the mean MAE and NMAE values across all stations for every available feature, and the bottom two images of Figure 7 visually compare the NMAE values with violin and density plots.
@@
Table 11: Mean and standard deviation of MAE and NMAE values across all stations. Feature names from left to right: t2m_max, t2m_mean, t2m_min, total_precipitation_mean, wind_speed_mean, wind_direction_at_max_speed, wind_speed_max, surface_pressure_max, and surface_pressure_min.
The results indicate that the temperature features are highly reliable. Even in the worst-performing station, the mean Mean Absolute Error (MAE) remains as low as $2.173$, with a standard deviation of $1.215$. Assuming a normal distribution of errors, this implies that in 95% of the cases, the discrepancy between AEMET measurements and IberFire data does not exceed 4°. Given that temperature values at a single station can vary by up to 35°during the year, this error value is relatively small and supports the robustness of the temperature variables in the IberFire datacube. Moreover, as the temperature values are derived from ERA5-Land, a dataset rigorously validated [ERA5_validation] by the Copernicus Climate Data Store, their reliability is further reinforced.
In comparison, the precipitation and surface pressure features exhibit higher NMAE values than temperature. Nevertheless, these variables are also sourced from Copernicus ERA5-Land without suffering much transformations. Consequently, despite their relatively higher error metrics, the integrity and reliability of these features remain supported by the quality and consistency of the underlying data source.
@@ 4.4 Fire risk mapping validation
To evaluate the practical applicability of the IberFire datacube in real-world scenarios, a fire risk mapping exercise was performed. The objective was to assess whether the datacube features can support the generation of reliable fire risk maps when used in combination with standard machine learning techniques.
@@
Forest fires are inherently rare events, resulting in a highly imbalanced distribution of the is_fire feature, with significantly fewer fire instances relative to non-fire occurrences. To address this imbalance during model training, a balanced dataset was retrieved from IberFire by including all fire instances recorded between 2008 and 2023, complemented by an approximately equal number of randomly selected non-fire instances from the same dates. Due to the stochastic nature of the sampling process of non-fire instances, the final class counts were not exactly equal. The resulting training dataset consisted of 140,399 instances, comprising 70,476 fire cases and 69,923 non-fire cases. For each selected instance corresponding to a unique spatio-temporal cell, all input features from the previous day (selecting the adequate CLC and popdens versions) were extracted from the IberFire datacube and stored in CSV format, and as target variable the is_fire of the selected instance was retrieved. This one-day difference between the input features and the output variable ensures that predictions can be made using information from the previous day.
@@
Following construction of the dataset, an XGBoost classifier was trained on the extracted instances. To reduce the risk of overfitting and ensure the generalisation capability of the model, a cross-validation strategy was applied. After model development, an independent test set was generated by retrieving a new balanced dataset, consisting of all available fire instances from the year 2024 along with a comparable number of randomly sampled non-fire instances from the same dates. The trained model was then evaluated on this test set, yielding an accuracy of 86%. Furthermore, the model was also tested on all the instances of 2024 (around $3.1\times 10^{9}$ instances) and achieved an Area Under the Receiver Operating Characteristic (AUROC) of 0.95, which indicates a high degree of predictive accuracy and robustness in real-world forecasting scenarios.
@@
In addition to this quantitative evaluation, the classifier was used to generate daily fire risk predictions for the entire year of 2024. For each day, all corresponding spatio-temporal instances were extracted from the IberFire datacube and converted into CSV format to serve as input for the model. Predictions were then computed for each instance. Subsequently, the auxiliary features x_index and y_index were used to construct the daily fire risk raster maps. Figure 8 presents a selection of predicted fire risk maps overlaid with the actual forest fire occurrences, enabling visual comparison between the predicted risk predictions and observed fires. Furthermore, the public repository associated with IberFire presents all the fire risk maps as an animation (see Section 6 for code availability).
Figure 8: Example of fire risk prediction maps. Left: predictions for 13/07/2024. Right: predictions for 20/12/2024.
@@
Visual inspection of the resulting fire risk maps revealed a high degree of spatial and temporal coherence and alignment with historical fire incidence patterns illustrated in Figure 1. Notably, areas with historically high fire activity, such as Galicia, Asturias, and regions along the Portuguese border, consistently exhibited higher predicted risk levels, reinforcing the model’s ability to capture meaningful spatial trends in fire susceptibility. Furthermore, as Figure 8 shows, the model is capable of making accurate predictions and assigning high risk to areas where forest fires actually occur. This is particularly evident in the winter predictions, which demonstrate the model’s ability to assign generally low risk across most regions while accurately identifying localised areas with a high likelihood of fire ignition.
@@ 5 Usage notes
The IberFire datacube was developed to support the modelling of wildfire occurrence risk across the Spanish territory, excluding the Canary Islands, Ceuta, and Melilla. A spatio-temporal structure was implemented to fulfil this objective, providing daily data from 01/12/2007 to 31/12/2024 over a regular grid of 1km $\times$ 1km spatial resolution. This design enables the training and validation of ML and DL models for wildfire risk prediction. The datacube is intended exclusively for mainland Spain and the Balearic Islands. To ensure that only relevant cells with complete and valid data are included in modelling workflows, users should filter the dataset accordingly using the variable is_spain (by selecting the instances where is_spain = 1).
@@
The datacube includes daily records beginning in December 2007; however, fire occurrence data were retrieved only from 01/01/2008 onward, and all cells for December 2007 were explicitly set as non-fire. This initial month is included to support the modelling of DL models that require a sequence of previous dates, like long short-term memory (LSTM) networks. Thus, even though fire predictions focus on the period from 2008 to 2024, the extended time frame ensures that models requiring temporal dependencies can be effectively trained.
@@
Prediction targets can be defined using either the is_fire variable, which indicates whether a fire occurred in a given cell on a specific date, or the is_near_fire variable, which flags cells located in proximity to a fire event. The use of the auxiliary variables is_spain, x_index and y_index as model inputs is not recommended. These features were introduced to facilitate data filtering, but they lack geospatial meaning beyond the cell grid. Their values are specific to the internal indexing of the datacube and are not transferable to other geographic regions. For modelling spatial geographical location, instead of x_index and y_index, the variables x_coordinate and y_coordinate, which represent the actual projected coordinates of each grid cell on the EPSG:3035 CRS, can be used as they retain geographic meaning and are suitable for location-aware modelling.
The remaining features in the datacube are suitable for use as explanatory variables in predictive models. Users can decide whether to integrate the baseline model FWI as an input feature, which when included, would result in a total of 120 different features for fire risk modelling.
@@
To incorporate information about the Spanish territorial division, it is recommended to apply one-hot encoding to the categorical AutonomousCommunities feature. This results in 17 binary features corresponding to the autonomous communities listed in Table 2, excluding the Canary Islands, Ceuta, and Melilla, which are not part of the study area. Similarly, a month feature can be derived from the temporal coordinate of each spatio-temporal cell. As different months exhibit distinct seasonal fire risk patterns, it is also advisable to apply one-hot encoding to this variable.
@@
It is important to note that some features, such as those derived from CLC and population density (popdens), are not static but also not temporally continuous. These variables are associated with a specific year of update and thus are included in the datacube with spatial dimensions only $(x,y)$, rather than full spatio-temporal indices $(x,y,t)$. When preprocessing the datacube for modelling applications or conversion to tabular formats such as CSV or DataFrame structures, it is crucial to ensure that for each spatio-temporal cell $(x,y,t)$, only information available at the time $t$ of the cell is used. For instance, when modelling a record for July 15, 2010, the appropriate population density value to use is popdens_2010, which should be assigned to a generic popdens column. Similarly, for a CLC-derived feature such as CLC_urban_fabric_proportion, the correct value to assign would be taken from CLC_2006_urban_fabric_proportion, since the next update in 2012 would not yet be available at that date. Careful consideration is required to prevent potential data leakage, particularly in the case of land cover features that might implicitly reflect fire occurrence. The CLC dataset contains a class representing the area that was burned, which, if taken from a dataset version updated after the date being modelled, could inadvertently introduce information correlated with the target variable (is_fire). This leakage could compromise the integrity and validity of predictive modelling results.
To create practically useful models, the instance extraction process should leverage historical data. For instance, the feature values from one day can be used to predict the value of the is_fire feature for the following day. For more sophisticated models, such as LSTM networks, a larger window of historical data can be used.
@@
Directly comparing the trained models with the baseline Fire Weather Index (FWI) is not straightforward, as AI classifiers typically produce probabilistic outputs, while the FWI generates continuous, regression-like values. To enable classification using the FWI, it is necessary to discretize its outputs. Based on the thresholds in Table 1, FWI values can be clipped at a maximum of 50 and linearly scaled to the [0, 1] range to approximate probabilities. Using this approach, cells with FWI values below 25 are classified as ‘non-fire’, while those above 25 are classified as ‘fire’, in accordance with the fire risk categories.
An essential final consideration when working with the IberFire datacube is the substantial computational resources required for data processing. Although the total disk size of the datacube is approximately 29GB, each spatio-temporal feature is stored in a highly compressed format. When decompressed, a single feature represented as a float32 array occupies around 25GB of memory, making it impractical to load multiple features into RAM simultaneously on standard hardware. The xarray package handles this problem by chunking the data. For lightweight tasks such as data visualization or exploratory analysis, standard hardware is sufficient. Similarly, retrieving a balanced CSV training dataset, including all fire instances of IberFire and an equal number of non-fire instances, can be done with standard hardware, as well as generating daily fire risk maps with all the daily instances. However, more demanding operations, such as the addition or generation of new spatio-temporal variables derived from combinations of existing features, require significantly more memory. Therefore, all processing steps involved in the construction of IberFire, including preprocessing and training dataset extraction, were executed on a machine equipped with 128GB of RAM.